Suppose you have 50.4 mL of a solution of H3PO4. You titrate it with a 0.10 M solution of LiOH and use 15.2 mL to reach the endpoint.

A.) What is the concentration of H3PO4?

B.) What was the pH of the original H3PO4 solution?
Answer #1

A)

Consider reaction, H3PO4 (aq) + 3 LiOH (aq) Li3PO4 (aq) + 3 H2O (l)

From reaction, Stoichiometric ratio = no. of moles of acid / no. of moles of base = 1 / 3

We have correlation, M acid V acid = M base V base Stoichiometric ratio

M acid = M base V base Stoichiometric ratio /  V acid

= 0.10 M 15.2 ml ( 1/3 ) / 50.4 ml

= 0.01005 M

ANSWER : [ H3PO4 ] = 0.01005 M

B)

Consider reaction, H3PO4 (aq) + H2O (l) H3O + (aq) + H2PO4- (aq)

Equilibrium constant for above reaction is , K a = [ H3O + ] [ H2PO4- ] / [ H3PO4 ] = 7.11 10 -03 .

Let's use ICE table to find out equilibrium concentrations of  H3O +, H2PO4- and H3PO4.

Concentration ( M ) H3PO4 H3O + H2PO4-
I 0.01005
C - X + X +X
E 0.01005 - X X X

Substituting these values in K a expression, we get

K a = ( X ) ( X ) / 0.01005 - X = 7.11 10 -03 .

X 2 / 0.01005 - X = 7.11 10 -03 .

X 2 = 0.01005 - X ( 7.11 10 -03 )

X 2 = 7.146 10 -05 - 7.11 10 -03 X

X 2 + 7.11 10 -03 X - 7.146 10 -05 = 0

Comparing to a X 2 + b X + c = 0 , we get a = 1 , b = 7.11 10 -03 and c = - 7.146 10 -05 .

Now, solve for X.

X = - b +/- b 2 -4 ac / 2 a

X = - b +/- ( 7.11 10 -03 ) 2 - 4 ( 1) ( - 7.146 10 -05 ) / 2 (1)

X = - b +/- 3.3637 10 -04 / 2

X = - b +/- 0.01834 / 2

X = -  7.11 10 -03 + 0.01834 /2 = 0.005615

or

X = -  7.11 10 -03 - 0.01834 /2 = - 0.012725

Acceptable value of X is 0.005615.

[ H3O + ] = [ H2PO4- ] = X = 0.005615 M

We have relation, pH = -log [ H3O + ]

pH = - log ( 0.005615 ) = 2.25

ANSWER : pH of H3PO4 solution is 2.25